Solve the determinant
3 x
1 1 = 2
(the 3 x and 1 1 are inside determinant signs)

Answer by venugopalramana(3286)

Solve the determinant
3 x
1 1 = 2
that is
3*1-1*x=2
3-x=2
x=3-2=1
x=1